# CompMarkGS: Robust Watermarking for Compressed 3D Gaussian Splatting

### v5

@@ D.3 Effectiveness of HSV loss
To address this, we introduce an HSV loss that aligns more closely with the perceptual characteristics of the human visual system. HSV color space separates chromatic components (Hue and Saturation) from luminance (Value), making it more suitable for detecting and reducing color artifacts. As shown in Fig. 6, binary masks of HSV loss successfully detect color artifacts in each color channel. Moreover, Fig. 7 shows that HSV loss allows 3DGS to focus its optimization on minimizing color artifacts, maintaining high bit accuracy. These results show that HSV loss effectively reduces color artifacts while preserving high bit accuracy and rendering quality.
@@
D.4 Robustness to model distortion comparison
@@
Fig.8 and Fig.9 show the bit accuracy under different levels of model distortion. We conduct three model distortions: 1) adding Gaussian noise to all model parameters, 2) randomly removing anchors, and 3) randomly cloning anchors. Across different distortion strengths, our proposed method consistently outperforms the baseline. Notably, under prune and clone distortion settings, our method exhibits less performance degradation, owing to the incorporation of both pruning and anchor growing during training in CompMarkGS. These results demonstrate that our method achieves superior performance over the baseline, even under severe model distortions.
Figure 8: Comparison of the baselines and our method under various model distortions and distortion strengths. We conduct three model distortions: 1) Gaussian noise, 2) cloning anchors, and 3) pruning anchors. The blue line represents the results of our method. Results represent the average score across Blender, LLFF, and Mip-NeRF 360 datasets using 48-bit messages. All baselines were tested within an anchor-based 3DGS framework, employing HAC.
Figure 9: Comparison of the baselines and our method under various model distortions and distortion strengths. We conduct three model distortions: 1) Gaussian noise, 2) cloning anchors, and 3) pruning anchors. The blue line represents the results of our method. Results represent the average score across Blender, LLFF, and Mip-NeRF 360 datasets using 48-bit messages. All baselines were tested within an anchor-based 3DGS framework, employing ContextGS.
D.5 Performance across compression levels
@@
Compression plays a crucial role in determining the trade-off between model size and performance in terms of bit accuracy and rendering quality. Fig.10 shows the relationship among model size, bit accuracy, and rendering quality under two different compression schemes: HAC [7] and ContextGS [39]. Fig.10 indicates that a lower compression level leads to better bit accuracy and rendering quality. Our method consistently outperforms other methods, achieving the highest bit accuracy and rendering quality across varying model sizes. Notably, for HAC compression, we achieve bit accuracy of $90\%$ with model size less than $10$ MB, while other methods only reach $60\%$ under the same model size. These results highlight the effectiveness and robustness of CompMarkGS.
@@
D.6 Additional qualitative results

### v6

@@ E.2 Effectiveness of HSV loss
When only RGB-based loss functions are used for watermark embedding, the optimization of 3DGS fails to consider the characteristics of the human visual system. As a result, 3DGS with a watermark renders color artifacts in the rendered images, particularly in regions with strong chromatic and luminance variations. These artifacts are observed in Fig.7 and Fig.8, despite achieving high bit accuracy. To address this, we introduce an HSV loss that aligns more closely with the perceptual characteristics of the human visual system. HSV color space separates chromatic components (Hue and Saturation) from luminance (Value), making it more suitable for detecting and reducing color artifacts. As shown in Fig. 7, binary masks of HSV loss successfully detect color artifacts in each color channel. Moreover, Fig. 8 shows that HSV loss allows 3DGS to focus its optimization on minimizing color artifacts, maintaining high bit accuracy. These results show that HSV loss effectively reduces color artifacts while preserving high bit accuracy and rendering quality.
@@
E.3 Robustness under various conditions
Robustness to pruning-based compression.
[…]
Method Bit Acc. (%) ↑ Size (MB) ↓ WateRF 92.65 / 67.29 248.97 / 84.77 GaussianMarker 92.89 / 66.67 418.66 / 141.84 3D-GSW 91.93 / 67.06 228.62 / 77.46 CompMarkGS 96.09 / 83.50 186.71 / 61.61
Robustness to general quantization.
[…]
Table 9: Robustness comparison against general quantization, showing the bit accuracy drop on the Mip-NeRF 360 dataset (48-bit messages) after quantizing models from float32 to int8.
Method Bit Acc. (% )↑ PSNR ↑ SSIM ↑ LPIPS ↓ WateRF 93.20 / 69.20 26.09 / 12.17 0.788 / 0.135 0.260 / 0.610 GaussianMarker 92.49 / 66.02 26.33 / 15.83 0.775 / 0.407 0.282 / 0.588 3D-GSW 91.08 / 66.20 18.53 / 14.93 0.667 / 0.374 0.322 / 0.662 CompMarkGS 95.59 / 90.11 26.56 / 21.73 0.777 / 0.700 0.248 / 0.336
E.4 Generality and Compatibility
Robustness to adversarial attacks.
[…]
Table 10: Robustness comparison against PGD adversarial attacks. Results are for 48-bit messages, showing bit accuracy before and after the attack targeting the HiDDeN decoder.
Method Bit Acc. (None)↑ Bit Acc. (PGD $\epsilon=0.1$) ↑ WateRF 92.01 56.06 GaussianMarker 87.95 57.53 3D-GSW 79.25 57.56 CompMarkGS 94.03 84.33
Decoder independence.
[…]
Table 11: Performance comparison with an alternative decoder (DINO-based SSL). Results are for 48-bit messages, averaged over three datasets, showing performance before and after HAC and ContextGS compression.
Methods Bit Accuracy (%) ↑ PSNR ↑ SSIM ↑ LPIPS ↓ HAC + WateRF 91.02 / 54.40 27.36 / 13.63 0.850 / 0.457 0.174 / 0.574 HAC + GaussianMarker 92.00 / 58.34 27.05 / 13.54 0.840 / 0.460 0.193 / 0.571 HAC + 3D-GSW 90.96 / 53.48 19.57 / 13.13 0.628 / 0.439 0.295 / 0.572 HAC + GuardSplat 79.77 / 52.72 16.29 / 12.28 0.584 / 0.425 0.417 / 0.609 HAC + CompMarkGS (SSL) 93.54 / 93.50 27.21 / 27.16 0.847 / 0.842 0.190 / 0.192 HAC + CompMarkGS (HiDDeN) 95.95 / 95.92 27.68 / 27.65 0.856 / 0.852 0.171 / 0.177 ContextGS + WateRF 92.01 / 90.36 26.64 / 26.47 0.843 / 0.832 0.183 / 0.185 ContextGS + GaussianMarker 91.24 / 87.95 26.89 / 26.54 0.839 / 0.827 0.195 / 0.200 ContextGS + 3D-GSW 88.28 / 79.75 19.50 / 19.83 0.627 / 0.617 0.294 / 0.299 ContextGS + GuardSplat 73.20 / 67.16 16.90 / 16.86 0.653 / 0.628 0.353 / 0.360 ContextGS + CompMarkGS (SSL) 93.09 / 93.00 25.86 / 25.82 0.829 / 0.828 0.220 / 0.221 ContextGS + CompMarkGS (HiDDeN) 94.36 / 94.03 27.60 / 27.55 0.845 / 0.844 0.172 / 0.173
E.5 Robustness to image distortion comparison
This section provides the detailed quantitative results for the experiment on robustness to image distortion, which is visually summarized in Fig.4 of the main paper. We evaluate watermark resilience by applying six post-processing distortions to the rendered images: Gaussian noise, rotation, scaling, Gaussian blur, crop, and JPEG compression. Tab.12 details the bit accuracy for each distortion condition, both before (left) and after (right) compression. As the results show, our proposed method, CompMarkGS, consistently exhibits superior robustness to existing methods across most distortion and compression scenarios.
Bit Accuracy(%) $\uparrow$ Methods No Distortion Gaussian Noise ($\sigma$ = 0.1) Rotation ($\pm\pi/6$) Scaling (75%) Gaussian Blur ($\sigma$ = 0.1) Crop (40%) JPEG Compression (50% quality) HAC + WateRF 91.02 / 54.40 89.75 / 55.90 84.92 / 56.16 86.17 / 55.38 89.33 / 56.34 87.00 / 55.30 84.00 / 57.47 HAC + GaussianMarker 92.00 / 58.34 59.08 / 58.33 58.75 / 58.33 58.33 / 58.33 59.08 / 58.33 61.42 / 58.17 58.25 / 58.33 HAC + 3D-GSW 90.96 / 53.48 89.50 / 52.50 86.67 / 52.25 86.42 / 52.08 86.67 / 52.25 86.92 / 52.50 83.83 / 53.25 HAC + GuardSplat 79.77 / 52.72 80.67 / 50.17 74.42 / 53.33 81.33 / 50.75 81.50 / 50.33 65.08 / 50.00 74.83 / 52.67 HAC + CompMarkGS 95.95 / 95.92 96.00 / 95.50 92.25 / 92.08 91.83 / 91.67 95.83 / 95.58 90.67 / 90.83 87.75 / 88.83 ContextGS + WateRF 92.01 / 90.36 92.25 / 91.50 88.75 / 88.00 89.58 / 88.83 92.75 / 91.50 89.50 / 88.25 86.41 / 84.58 ContextGS  + GaussianMarker 91.24 / 87.95 59.08 / 58.92 59.00 / 58.75 58.25 / 58.17 59.17 / 58.92 61.08 / 60.67 58.25 / 58.25 ContextGS  + 3D-GSW 88.28 / 79.25 87.92 / 80.67 85.08 / 78.75 84.42 / 77.08 88.00 / 81.42 84.33 / 78.50 81.58 / 72.58 ContextGS  + GuardSplat 73.20 / 67.16 73.92 / 65.83 67.00 / 61.83 74.25 / 66.67 74.33 / 67.08 62.33 / 54.75 63.75 / 61.67 ContextGS + CompMarkGS 94.36 / 94.03 93.04 / 92.98 90.17 / 90.33 90.46 / 90.06 92.96 / 92.81 89.63 / 89.29 85.21 / 85.31
E.6 Robustness to model distortion comparison
@@
Fig.9(a) and Fig.9(b) show the bit accuracy under different levels of model distortion. We conduct three model distortions: 1) adding Gaussian noise to all model parameters, 2) randomly removing anchors, and 3) randomly cloning anchors. Across different distortion strengths, our proposed method consistently outperforms the baseline. Notably, under prune and clone distortion settings, our method exhibits less performance degradation, owing to the incorporation of both pruning and anchor growing during training in CompMarkGS. These results demonstrate that our method achieves superior performance over the baseline, even under severe model distortions.
(a) HAC compression.
(b) ContextGS compression.
Figure 9: Comparison of baselines and our method under Gaussian noise, cloning, and pruning. Results are averaged over Blender, LLFF, and Mip-NeRF 360 with 48-bit messages.
Table 13: Robustness to model distortion before (left) and after (right) compression.
Bit Accuracy(%) $\uparrow$ Methods No Distortion Gaussian Noise ($\sigma$ = 0.005) Clone (50%) Prune (20%) HAC + WateRF 91.02 / 54.40 81.03 / 53.38 89.78 / 54.04 87.33 / 53.70 HAC + GaussianMarker 92.00 / 58.34 73.18 / 58.35 90.23 / 58.29 87.90 / 87.89 HAC + 3D-GSW 90.96 / 53.48 80.17 / 53.11 89.62 / 53.17 86.80 / 53.18 HAC + GuardSplat 79.77 / 52.72 67.32 / 51.21 76.12 / 56.28 71.25 / 59.13 HAC + CompMarkGS 95.95 / 95.92 82.13 / 77.13 95.33 / 95.09 94.67 / 94.25 ContextGS + WateRF 92.01 / 90.36 82.36 / 79.91 90.70 / 88.95 88.29 / 86.20 ContextGS  + GaussianMarker 91.24 / 87.95 72.84 / 71.78 89.43 / 86.05 86.97 / 83.56 ContextGS  + 3D-GSW 88.28 / 79.75 76.94 / 68.55 86.89 / 77.99 82.87 / 72.70 ContextGS  + GuardSplat 73.20 / 67.16 63.24 / 59.84 71.39 / 70.11 67.12 / 65.19 ContextGS + CompMarkGS 94.36 / 94.03 87.03 / 83.08 93.37 / 93.23 92.38 / 92.57
E.7 Performance across compression levels
@@
Compression plays a crucial role in determining the trade-off between model size and performance in terms of bit accuracy and rendering quality. Fig.10 shows the relationship among model size, bit accuracy, and rendering quality under two different compression schemes: HAC (Chen et al., 2024a) and ContextGS (Wang et al., 2024a). Fig.10 indicates that a lower compression level leads to better bit accuracy and rendering quality. Our method consistently outperforms other methods, achieving the highest bit accuracy and rendering quality across varying model sizes. Notably, for HAC compression, we achieve bit accuracy of $90\%$ with model size less than $10$ MB, while other methods only reach $60\%$ under the same model size. These results highlight the effectiveness and robustness of CompMarkGS.
(a) HAC compression.
(b) ContextGS compression.
@@
Appendix F Computing resources
We evaluate the rendering performance of our method and baselines (Huang et al., 2025; Jang et al., 2024; 2025) by measuring FPS. As shown in Tab.14, our method exceeds the real-time rendering threshold of $30$ FPS across the Blender (Mildenhall et al., 2021), LLFF (Mildenhall et al., 2019), and Mip-NeRF 360 (Barron et al., 2022) datasets. These results demonstrate the practical applicability and rendering efficiency of our proposed CompMarkGS.
FPS ↑ Methods Blender LLFF Mip-NeRF 360 WateRF 199.30 19.31 41.98 GaussianMarker 166.45 16.24 28.54 3D-GSW 220.60 22.67 50.64 GuardSplat 207.56 21.93 41.28 CompMarkGS 230.39 35.30 67.58
Appendix G Additional qualitative results